Interest rate. In 1972, Bob purchased a new Datsun 240Z for $3,000. Datsun later changed its name to Nissan, and the 1972 Datsun 240Z became a classic. Bob kept his car in excellent condition and in 2002 could sell the car for five times what he originally paid. What was Bobs annualized rate of return for the 30 years he owned this car? If he keeps the car for another thirty years and earns the same rate, what could he sell the car for in 2032? What was Bobs annualized rate of return for the 30 years he owned this car? (Round to two decimal places.)
Answer #1
Original price 3000
Projected sale price 15000
Time 30 years
15000 3000*(1+r)^30
5 (1+r)^30
1+r= 5^(1/30)
1+r= 1.055113064
r= 0.055113064
r= 5.511%
Expected price after next 30 years 15000*(1+5.511%)^30
Expected price after next 30 years                           75,000.00